# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v2

@@ Appendix D Experimental Details and Additional Results
Apps. D.1-D.4 provide additional details regarding the data and GPHLVM training process for each of the experiments described in Sec. 5. App D.5 provides additional insights into the hand grasps experiment of Sec. 5.4.
@@ D.1 C-shape
The proof-of-concept experiment on the $\mathsf{C}$-shape dataset was designed as a minimalist example with a simple setup. The training data consists of $N=1000$ two-dimensional points arranged in a $\mathsf{C}$ shape and scaled to fit within the unit circle. In the Euclidean case, these datapoints are used as both the latent points and the observations, so that $\bm{X}=\bm{Y}\in\mathbb{R}^{1000\times 2}$. In the hyperbolic case, the datapoints can be interpreted as elements of the Poincaré model. We represent each latent point and observation in the Lorentz model using the Poincaré to Lorentz mapping $\bm{x}_{\mathcal{L}}=\frac{1}{1-\|\bm{x}\|^{2}}\begin{bmatrix}1+\|\bm{x}\|^{2}&2\bm{x}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}\in\mathbb{H}^{2}_{\mathcal{L}}$. In this proof-of-concept experiment, we manually set the kernel variance, lengthscale, and noise variance of both GPLVM and GPHLVM to $(\tau,\kappa,\sigma^{2}_{y})=(0.7,0.15,0.69)$. Additionally, for the GPHLVM, we used the 2D hyperbolic SE kernel with $L=3000$ rejection samples. We used constant zero mean for both models.
The start- and end-points of the geodesics shown in Fig. 2 correspond to one of the $\mathsf{C}$-shape latent points. The geodesics are represented by $M=25$ regularly-spaced points. Geodesics on the base Euclidean manifold correspond to straight lines from $\bm{x}_{0}$ to $\bm{x}_{M-1}$. In the hyperbolic case, the geodesics on the base manifold are given as $\bm{x}_{i}=\text{Exp}_{\bm{x}_{0}}(t_{i}\text{Log}_{\bm{x}_{0}}(\bm{x}_{M-1}))$ with $t_{i}=i/(M-1)$ and $i=\{0,...,M-1\}$. The geodesics on the base manifolds were used as initialization for the pullback geodesic optimization. To compute the pullback geodesic, we applied Riemannian Adam for $200$ steps with a learning rate of $0.005$, optimizing the combined loss function (24) based on both curve energy and spline energy, as introduced in Sec. 3.3. In this experiment, the curve energy and spline energy were weighted equally, i.e., $\lambda=1$ in Eq. (23).
Finally, we generated a $110\times 110$ grid of points to visualize the volume of the pullback metric, filtering out those points lying outside the unit circle in the hyperbolic case. Then, we evaluated the pullback metric tensor at each remaining grid point. In the Euclidean case, we computed the volume at each grid point $\bm{x}^{*}$ by evaluating the determinant of the $2\times 2$ pullback metric tensor, $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}^{*}})}$. In the hyperbolic case, the pullback metric tensor is a $3\times 3$ matrix that lies on the $2$-dimensional tangent space, leading to one of its eigenvalues being zero. To address this, we compute the volume as the product of the two nonzero eigenvalues.
@@ D.2 MNIST Digits
We use a subset of the MNIST dataset composed of $100$ datapoints for each of the classes 00, $1$, $2$, $3$, $6$, and $9$. Each datapoint is represented as a $28\times 28$ binary image. We optimize the parameters of the GPHLVM using Riemannian Adam for $500$ steps, with a learning rate of $0.05$. We set a Gamma prior with concentration $\alpha=2$ and rate $\beta=2$ on the kernel lengthscale and a Gamma prior with concentration $\alpha=5$ and rate $\beta=0.8$ on the kernel variance. The embeddings are initialized using PCA.
D.3 Multi-cellular Robot Design
We consider multi-cellular robots composed of a $5\times 5$ grid of cells. Possible cell types are empty, solid, soft, a horizontal actuator or a vertical actuator. We generate a dataset of $216$ multicellular robots following the hierarchical approach of Dong et al. (2024), which we explain next for completeness. The root nodes of the hierarchy are two coarse robots composed of only vertical and only horizontal actuator cells. A component is defined as a group of cells of the same type, i.e., the root robots have a single component. Children nodes of each robot are designed by dividing large components into $2$ smaller components using K-means and by assigning a different type of cell to one of these components. This process is repeated iteratively to obtain robots with up to $8$ components. This results in a tree of multi-cellular robots, with coarse-grained robots close to the roots and fine-grained robots at the leaves.
To account for the hierarchical structure associated with the tree structure of the multi-cellular robots, we incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al. (2024). This is achieved by leveraging the stress loss,
@@ D.3 Multi-cellular Robot Design
to match the pairwise latent distances with the tree distances, and where $c_{i}$ denotes the tree node, $\operatorname{dist}_{\mathbb{G}}$, and $\operatorname{dist}_{\mathcal{M}}$ are the graph distance and the geodesic distance on $\mathcal{M}$, respectively. Therefore, training the GPHLVM is achieved by solving,
@@
where $\gamma$ is a parameter balancing the two losses and ¯$\bar{\ell}_{\text{stress}}(\bm{X})$ is the stress loss averaged over the embeddings. We set $\gamma=6000$. Moreover, we use a Gamma prior with concentration $\alpha=2$ and rate $\beta=2$ on the kernel lengthscale. We optimize the parameters of the GPHLVM using Riemannian Adam for $500$ steps, with a learning rate of $0.05$.
@@ D.4 Hand Grasps Generation
For the hand grasp experiment, we used data from the KIT Whole-Body Motion Database (Mandery et al., 2016). The dataset consists of $38$ trajectories, with motion recordings from two human subjects (IDs $2122$ and $2123$). Each subject performed $19$ different grasp types [^1] The three-finger sphere grasp of the GRASP taxonomy (Feix et al., 2016) is missing from the dataset. of the GRASP taxonomy (Feix et al., 2016). Each grasp consists of the subject grasping an object from a table, lifting it, and placing it back.
We applied several preprocessing steps to the recorded data before training the GPHLVM model: (1) We applied a low-pass filter to remove high-frequency noise; (2) We trimmed the start and end of each trajectory to focus solely on the motion from the initial resting pose to the point where the grasp was completed. Since detecting the exact point of grasp completion is non-trivial, we cut off the trajectories at the moment when the grasped object was first moved by the subject; (3) We subsample the trajectories; and (4) We centered the data to allow for the use of a zero mean function in the GPHLVM model. After preprocessing, each trajectory is composed of $30$ to $40$ data points. Stacking all trajectories together results in a total of $N=1321$ data points $\bm{Y}\in\mathbb{R}^{1321\times 24}$.
In contrast to the previous experiments, this experiment involves trajectory data. To preserve the trajectory structure during training, we incorporated a dynamics prior $p(\bm{X}_{2:N}\mid\bm{X}_{1:N-1})$, similar to Wang et al. (2008), but using the wrapped Gaussian distribution (38). We also used back constraints $\bm{X}=k(\bm{Y},\bm{Y})\bm{C}$, which allow for a smooth mapping from the observation space to the latent space. The back-constraints kernel is defined as a Euclidean SE kernel with variance $\tau=1$ and a lengthscale $\kappa=0.4$. Moreover, each grasp type is identified with a lead node of the quantitative taxonomy of hand grasps (Stival et al., 2019). More specifically, the first point of each trajectory, which corresponds to the resting state, is assigned to the root node $c_{0}$ of the taxonomy graph, and the last point to the taxonomy node representing the corresponding grasp type. The taxonomy node $c_{14}$, for instance, represents the stick grasp. The number of edges between different grasp types in the taxonomy graph defines a distance function, $\text{d}_{\mathbb{G}}(c_{i},c_{j})$. To account for the hierarchical structure associated with the taxonomy, we additionally incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al. (2024). This is achieved by leveraging the stress loss (67) to match the pairwise latent distances with the graph distances. During training, we minimize the stress loss over the first and last latent points of each trajectory along with maximizing the GPHLVM marginal likelihood by optimizing,
@@
(a) GPLVM, $\mathbb{R}^{2}$
(b) GPLVM, $\mathbb{R}^{3}$
Figure 8: Hand grasps transition from a ring ( ) to a spherical grasp ( ) using Euclidean ( ) and pullback ( ) geodesics in (a) $\mathbb{R}^{2}$ and (b) $\mathbb{R}^{3}$. Top left: Latent embeddings. The latent trajectory points are colored according to their corresponding grasp class. The background color in $\mathbb{R}^{2}$ represents the pullback metric volume. Top right: Time-series plots of $2$ dimensions of the joint space. The top and bottom rows show the mean of the decoded hyperbolic ( ) and pullback geodesics ( ) with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded hyperbolic (top) and pullback (bottom) geodesics.
D.5 Additional Hand Grasps Results
This section provides further insights into the hand grasps experiment described in Sec. 5.4. First, we present additional visualizations of the decoded geodesics. While Fig. 5 only showed $2$ dimensions of the joint space, Fig. 7 depicts the decoded geodesics across all $24$ dimensions, i.e., over the $24$ degrees of freedom (DoF) of the hand model introduced in Mandery et al. (2016). This model counts two joints for the wrist, four joints for each finger, and two additional joints for the ring and little fingers to allow a better fist closure.
Second, we compare the GPHLVM and geodesics generated in Sec. 5.4 to their counterpart Euclidean models. We train two GPLVMs augmented with a taxonomy graph prior and a dynamic prior similarly as the GPHLVMs of Sec. 5.4. Fig. 8 show the resulting 2D and 3D latent spaces, alongside the Euclidean and Euclidean pullback geodesic interpolating from a ring to a spherical grasps and their corresponding decoded hand motions.
Third, we discuss the influence of initialization on optimizing the pullback geodesics. As described in Sec. D.4, we initialize the pullback geodesic with two concatenated hyperbolic geodesics: One from the ring grasp to the origin and another from the origin to the spherical grasp. This initialization incorporates knowledge about the expected path in the taxonomy graph. In contrast, Fig. 9 shows the optimized pullback geodesics when the optimization is initialized solely with a hyperbolic geodesic from the ring grasp to the spherical grasp. In the two-dimensional case, we observe that the optimized pullback geodesic path differs significantly from the one shown in Fig. 5. This discrepancy is likely due to the geodesic optimization getting stuck in a local minimum, which may be a consequence of the Monte Carlo approximation of the 2D hyperbolic SE kernel. Interestingly, the three-dimensional pullback geodesic does not exhibit this issue, further supporting the hypothesis that the Monte Carlo approximation may be responsible since the 3D case does not rely on this approximation. In the 3D case, the optimized pullback geodesic follows a path very similar to the one shown in Fig. 5.

### v3

@@ Appendix D Experimental Details and Additional Results
Apps. D.1-D.4 provide additional details regarding the data and GPHLVM training process for each of the experiments described in Sec. 5. App. D.5 provides additional insights into the hand grasps experiment of Sec. 4, and App. D.6 provides runtimes for the computation of pullback metrics and geodesics.
@@ D.1 C-shape
The proof-of-concept experiment on the $\mathsf{C}$-shape dataset was designed as a minimalist example with a simple setup to visualize and compare Euclidean and hyperbolic pullback metrics. The training data consists of $N=1000$ two-dimensional points arranged in a $\mathsf{C}$ shape and scaled to fit within the unit circle. In the Euclidean case, the datapoints are used as both latent points and observations, i.e., $\bm{X}=\bm{Y}\in\mathbb{R}^{1000\times 2}$. In the hyperbolic case, the datapoints can be interpreted as elements of the Poincaré model. We represent each latent point and observation in the Lorentz model using the Poincaré to Lorentz mapping $\bm{x}_{\mathcal{L}}=\frac{1}{1-\|\bm{x}\|^{2}}\begin{bmatrix}1+\|\bm{x}\|^{2}&2\bm{x}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}\in\mathbb{H}^{2}_{\mathcal{L}}$. We fully specify the LVMs by manually setting the kernel variance, lengthscale, and noise variance of both GPLVM and GPHLVM to $(\tau,\kappa,\sigma^{2}_{y})=(0.7,0.15,0.69)$. We used constant zero mean for both models. For the GPHLVM, we used the 2D hyperbolic SE kernel with $L=3000$ rejection samples.
The start- and end-points of the geodesics shown in Fig. 2 correspond to one of the $\mathsf{C}$-shape latent points. The geodesics are represented by $M=25$ regularly-spaced points. Geodesics on the Euclidean base manifold correspond to straight lines from $\bm{x}_{0}$ to $\bm{x}_{M-1}$. In the hyperbolic case, the geodesics on the base manifold are given as $\bm{x}_{i}=\text{Exp}_{\bm{x}_{0}}(t_{i}\text{Log}_{\bm{x}_{0}}(\bm{x}_{M-1}))$ with $t_{i}=i/(M-1)$ and $i=\{0,...,M-1\}$. The geodesics on the base manifolds were used as initialization for the pullback geodesic optimization. We obtain the pullback geodesics by minimizing the combined loss function (18) based on both curve energy and spline energy (see Sec. 3.3). We used Riemannian Adam for $200$ steps with a learning rate of $0.005$. In this experiment, the curve energy and spline energy were weighted equally, i.e., $\lambda=1$ in Eq. (17).
We visualize the volume of the pullback metrics on a $110\times 110$ grid of points, filtering out points lying outside the unit circle in the hyperbolic case. In the Euclidean case, we computed the volume at each grid point $\bm{x}^{*}$ by evaluating the determinant $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}^{*}})}$ of the pullback metric tensor $\bm{G}^{\text{P},\mathbb{R}}$. In the hyperbolic case, the pullback metric tensor is a $3\times 3$ matrix that lies on the $2$-dimensional tangent space, leading to one of its eigenvalues being zero. To address this, we compute the volume $\sqrt{\text{det}(\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}})}$ as the product of the two nonzero eigenvalues.
@@ D.2 MNIST Digits
We use a subset of the MNIST dataset composed of $100$ datapoints for each of the classes 00, $1$, $2$, $3$, $6$, and $9$. Each datapoint is represented as a $28\times 28$ binary image. We optimize the parameters of the GPHLVM using Riemannian Adam for $500$ steps with a learning rate of $0.05$. We set a Gamma prior with concentration $\alpha=2$ and rate $\beta=2$ on the kernel lengthscale and a Gamma prior with concentration $\alpha=5$ and rate $\beta=0.8$ on the kernel variance. The hyperbolic embeddings are initialized in the tangent space of the origin $\bm{\mu}_{0}$ using PCA and projected onto the manifold using the exponential map. Similarly, we optimize the parameters of the GPLVM using Adam for $500$ steps with a learning rate of $0.01$. The Euclidean embeddings are initialized using PCA.
D.3 Multi–cellular Robot Design
We consider multi–cellular robots composed of a $5\times 5$ grid of cells. Possible cell types are empty, solid, soft, a horizontal actuator or a vertical actuator. We generate a dataset of $216$ multi-cellular robots following the hierarchical approach of [16], which we explain next for completeness. The root nodes of the hierarchy are two coarse robots composed of only vertical and only horizontal actuator cells. A component is defined as a group of cells of the same type, i.e., the root robots have a single component. Children nodes of each robot are designed by dividing large components into $2$ smaller components using K-means and by assigning a different type of cell to one of these components. This process is repeated iteratively to obtain robots with up to $8$ components. This results in a tree of multi-cellular robots, with coarse-grained robots close to the roots and fine-grained robots at the leaves.
To account for the hierarchical structure associated with the tree structure of the multi-cellular robots, we incorporate graph-based priors in both GPHLVM and GPLVM as proposed in [25]. This is achieved by leveraging the stress loss,
@@ D.3 Multi–cellular Robot Design
to match the pairwise latent distances with the tree distances, and where $c_{i}$ denotes the tree node, $\operatorname{dist}_{\mathbb{G}}$, and $\operatorname{dist}_{\mathcal{M}}$ are the graph distance and the geodesic distance on $\mathcal{M}$, respectively. Therefore, training the GPHLVM and GPLVM is achieved by solving,
@@
where $\gamma$ is a parameter balancing the two losses and ¯$\bar{\ell}_{\text{stress}}(\bm{X})$ is the stress loss averaged over the embeddings. We set $\gamma=6000$. For the GPHLVM, we use a Gamma prior with concentration $\alpha=2$ and rate $\beta=2$ on the kernel lengthscale. We optimize the parameters of the GPHLVM using Riemannian Adam for $500$ steps with a learning rate of $0.05$ and the parameters of the GPLVM using Adam for $500$ steps with a learning rate of $0.01$.
@@ D.4 Hand Grasps Generation
For the hand grasp experiment, we used data from the KIT Whole-Body Motion Database [36]. The dataset consists of $38$ trajectories, with motion recordings from two human subjects (IDs $2122$ and $2123$). Each subject performed $19$ different grasp types [^1] The three-finger sphere grasp of the GRASP taxonomy [18] is missing from the dataset. of the GRASP taxonomy [18]. Each grasp consists of the subject grasping an object from a table, lifting it, and placing it back.
We applied several preprocessing steps to the recorded data before training the GPHLVM: (1) We applied a low-pass filter to remove high-frequency noise; (2) We trimmed the start and end of each trajectory to focus solely on the motion from the initial resting pose to the point where the grasp was completed. Since detecting the exact point of grasp completion is non-trivial, we cut off the trajectories at the moment when the grasped object was first moved by the subject; (3) We subsample the trajectories; and (4) We centered the data to allow for the use of a zero mean function in the GPHLVM. After preprocessing, each trajectory is composed of $30$ to $40$ data points. Stacking all trajectories together results in a total of $N=1321$ data points $\bm{Y}\in\mathbb{R}^{1321\times 24}$.
In contrast to the previous experiments, this experiment involves trajectory data. To preserve the trajectory structure during training, we incorporated a dynamics prior $p(\bm{X}_{2:N}\mid\bm{X}_{1:N-1})$, similar to [52], but using the wrapped Gaussian distribution (27). We also used back constraints $\bm{X}=k(\bm{Y},\bm{Y})\bm{C}$, which allow for a smooth mapping from the observation space to the latent space. The back-constraints kernel is defined as a Euclidean SE kernel with variance $\tau=1$ and a lengthscale $\kappa=0.4$. Moreover, each grasp type is identified with a lead node of the quantitative taxonomy of hand grasps [50]. More specifically, the first point of each trajectory, which corresponds to the resting state, is assigned to the root node $c_{0}$ of the taxonomy graph, and the last point to the taxonomy node representing the corresponding grasp type. The taxonomy node $c_{14}$, for instance, represents the stick grasp. The number of edges between different grasp types in the taxonomy graph defines a distance function, $\text{d}_{\mathbb{G}}(c_{i},c_{j})$. To account for the hierarchical structure associated with the taxonomy, we additionally incorporate graph-based priors in our GPHLVM as proposed by Jaquier et al. [25]. This is achieved by leveraging the stress loss (54) to match the pairwise latent distances with the graph distances. During training, we minimize the stress loss over the first and last latent points of each trajectory along with maximizing the GPHLVM marginal likelihood by optimizing,
@@
Refer to caption Refer to caption Refer to caption
(a) GPHLVM, $\mathbb{H}^{3}_{\mathcal{L}}$
Refer to caption Refer to caption Refer to caption
(b) GPLVM, $\mathbb{R}^{3}$
Figure 7: Top left: Embeddings of hand grasps colored according to their corresponding grasp class. The background color represents the pullback metric volume. The base manifold ( ) and pullback ( ) geodesics correspond to a transition from a ring ( ) to a spherical grasp ( ). Top right: Time-series plots of $2$ dimensions of the joint space showing the mean of the decoded geodesics with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded geodesics.
(a) GPHLVM, $\mathbb{H}^{2}_{\mathcal{L}}$
(b) GPLVM, $\mathbb{R}^{2}$
(c) GPHLVM, $\mathbb{H}^{3}_{\mathcal{L}}$
(d) GPLVM, $\mathbb{R}^{3}$
Figure 8: Pairwise error matrices between geodesic and taxonomy graph distances.
D.5 Additional Hand Grasps Results
This section provides further insights into the hand grasps experiment described in Sec. 4. First, Fig. 7 shows the 3D hyperbolic and Euclidean latent spaces obtained from the GPHLVM and GPLVM of Section 5, alongside the base manifold and pullback geodesic interpolating from a ring to a spherical grasps and their corresponding decoded hand motions.
Second, we discuss the stress values provided in Table 2, which measure the average difference between the hyperbolic distances between embeddings and the corresponding taxonomy graph distances. Figure 8 complements the statistics presented in Table 2 and depicts the pairwise error matrices between the graph and manifold distances for the 2D and 3D GPHLVMs and GPLVM. The left and top bar colors represent the embedding’s taxonomy classes following the color code of Figs. 5 and 7. The rest pose class (root node of the tree) is represented in white. We observe that, for all models, the pairwise errors between two leaf nodes (i.e., grasps), represented in the bottom-right corner of the plots, are higher than the pairwise errors between rest poses and grasps (i.e., root and leaf nodes), represented in the top-left, top-right, and bottom-left corners. This explains the large variance in the stress values of Table 2, and is due to the fact that a low error between the root and the leaves simply involves the leaf embeddings to be at the right distance from the origin, where the root is roughly located. Low errors between leaf nodes additionally involve the adequate relative placement of the corresponding embeddings. It is important to emphasize that the hyperbolic models lead to lower pairwise errors compared to their Euclidean counterparts for all taxonomy classes, as shown by the darker colors of Figs. 8(a) and 8(c) compared to Figs. 8(b) and 8(d). Moreover, the 3D hyperbolic model leads to the lowest pairwise errors overall.
Third, we present additional visualizations of the decoded geodesics. While Figs. 5 and 7 only show $2$ dimensions of the joint space, Fig. 9 depicts the decoded geodesics across all $24$ dimensions, i.e., over the $24$ degrees of freedom (DoF) of the hand model introduced in [36]. This model counts two joints for the wrist, four joints for each finger, and two additional joints for the ring and little fingers to allow a better fist closure.
@@ D.5 Additional Hand Grasps Results
Finally, we discuss the influence of initialization on optimizing the pullback geodesics. As described in Sec. D.4, we initialize the pullback geodesic with two concatenated hyperbolic geodesics: One from the ring grasp to the origin and another from the origin to the spherical grasp. This initialization incorporates knowledge about the expected path in the taxonomy graph. In contrast, Fig. 10 shows the optimized pullback geodesics when the optimization is initialized solely with a hyperbolic geodesic from the ring grasp to the spherical grasp. In the two-dimensional case, we observe that the optimized pullback geodesic path differs significantly from the one shown in Fig. 5. This discrepancy is likely due to the geodesic optimization getting stuck in a local minimum, which may be a consequence of the Monte Carlo approximation of the 2D hyperbolic SE kernel. Interestingly, the three-dimensional pullback geodesic does not exhibit this issue, further supporting the hypothesis that the Monte Carlo approximation may be responsible since the 3D case does not rely on this approximation. In the 3D case, the optimized pullback geodesic follows a path very similar to the one shown in Fig. 5.
@@
D.6 Runtimes
Table 3: Average computation times in seconds over $10$ computations of the pullback metric at a random point $\bm{x}$ and of a pullback geodesic between two random points $\bm{x},\bm{y}$.
Model Pullback metric Geodesic
GPLVM $\mathbb{R}^{2}$ $0.65\times 10^{-3}\pm 0.06\times 10^{-3}$ $4.14\pm 0.27$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ $44.2\times 10^{-3}\pm 1.3\times 10^{-3}$ $38.08\pm 0.52$
GPLVM $\mathbb{R}^{3}$ $0.63\times 10^{-3}\pm 0.01\times 10^{-3}$ $4.19\pm 0.23$
GPHLVM $\mathbb{H}^{3}_{\mathcal{L}}$ $1.77\times 10^{-3}\pm 0.07\times 10^{-3}$ $13.67\pm 0.33$
Table 3 shows the runtime measurements for the computation of pullback metrics and pullback geodesics of GPHLVM and GPLVM of the hand grasps generation experiment. The implementations are fully developed on Python, and the runtime measurements were taken using a Macbook Pro with Apple M3 Max chip with $14$-core CPU and $36$GB RAM. The main computational burden arises in the GPHLVM with 2D latent space, which is in contrast with the experiments using a 3D latent space. We attribute this increase of computational cost mainly to the 2D hyperbolic kernel, as also observed in [25]. This may be alleviated by reducing the number of samples or via more efficient sampling strategies, as briefly discussed in Section 6.